INTERNATIONAL COLLOQUIUM ON COMBINATORICS AND GRAPH THEORY a satellite conference of ECM2 Balatonlelle, Hungary, July 15-19, 1996

REGISTRATION FORM

Name: Sex: Affiliation:

Mailing address:(if different)

e-mail address: Accompanying person(s):

Do you intend to give a 20-minute contributed talk?

Yes No

Title:

Key words: I enclose my abstract

Yes No

Do you want us to make reservation for your accommodation?

Yes No

Date of arrival: Date of departure: If yes, what type of accommodation do you choose? Hotel Giuseppe

Single Double Appartment (4 pers.)

With whom would you like to share your room? I enclose the registration fee I will send the registration fee by ... Please return this form by May 15, 1996 to: J. BOLYAI MATHEMATICAL SOCIETY Budapest, Fő u. 68., H-1027, Hungary, or to the e-mail address comb96@math-inst.hu Date:

Signature